### v3

# Multiphoton Spectroscopy of a Dynamical Axion Insulator

Abstract
The unusual magnetoelectric transport present in Weyl semimetals and 3D topological insula- tors can be compactly understood as manifestations of a background axion field, which itself is determined by the microscopic band structure. In the presence of correlations, an additional axion quasiparticle may emerge as the collective excitations on top of the mean background field. Such modes couple nonlinearly to electric and magnetic fields, giving rise to a dynamical magnetoelectric response. However, unambiguous identification of this collective axion mode is challenging due to its inherent nonlinear dynamics. Here, we propose an all-optical protocol that utilizes a pump-probe setup for verifying and characterizing the transient dynamics of axion fields in three-dimensional insulator systems. In particular, we show that nonlinear Raman processes induce dynamical oscillations of the axion field that depend on the geometry of the incident electromagnetic fields. These oscillations manifest in the polarization and magnetization of the material, hence, can be subsequently measured using time-resolved Kerr rotation spectroscopy. Our results open a pathway towards using multi-photon and quantum pair spectroscopies to identify new correlated phases of quantum matter.
preprint: APS/123-QED
Introduction—Axionic particles were originally proposed in high-energy physics in order to solve the charge-parity problem as they were associated to the Nambu-Goldstone bosons of a new global axial U(1) symmetry [1, 2]. Today, axions are a leading dark matter candidate and could help explain the matter-antimatter asymmetry present in the universe [3]. While the axion has yet to be discovered in particle physics, its condensed matter counterpart has been theorized to exist in certain three-dimensional materials [4, 5, 6, 7, 8, 9, 10]. A characteristic identifier of the presence of an axion is a signature magnetoelectric response, leading to a modified form of Maxwell’s equations known as axion electrodynamics [11, 12, 13]; these give rise to physical phenomena which can broadly be classified as either transport or density responses. Axion-induced transport includes anomalous Hall currents which are related to the gradient of the axion, and a magnetic conductivity that relates to the axion’s first time derivative. Density responses have to do with the topological magnetoelectric effect where the axion acts as the coupling constant between magnetic field and polarization; and similarly between electric field and magnetization.
In correlated systems, the axion field has an associated collective excitation which couples nonlinearly to the electromagnetic field; this induces a dynamical magnetoelectric effect on top of the mean field response due to the background axion field [14, 15, 16, 17]. A variety of candidate materials have been put forward where axion field effects manifest in the electronic band structure. For example, axion fields naturally appear in correlated Weyl semimetals that undergo lattice translation symmetry breaking due to the emergence of charge density wave (CDW) order. In this case, the mean background axion field is determined by the separation of Weyl points in momentum space and in energy; while its dynamical contribution is associated with collective excitations of the phase of the CDW, i.e., the Nambu-Goldstone bosons [18, 19]. In the presence of uncompensated carrier densities the softening of the axionic collective mode may even lead to spatially inhomogeneous correlated phases [20]. 3D topological insulators (TI) can also support a static, background axion field which, due to time-reversal symmetry, acts as a $\mathbb{Z}_{2}$ topological invariant and takes quantized values $\theta=0$ or $\pi$. When interactions are turned on that spontaneously break time-reversal symmetry, as in the case of antiferromagnetic (AFM) order, $\theta$ is promoted to a dynamical quantity that takes continuous values between 0 and $2\pi$ [5]. Here, the fluctuations of the axion field are induced by the spin-wave excitations of the AFM order parameter [5, 21, 22].
The emergence of axion physics in condensed matter systems has ignited hope of studying such elusive physics in a tangible, tabletop experimental setting. While the high energy axion continues to go undetected, measuring axion quasiparticles in condensed matter systems have posed its own theoretical and experimental challenges due its inherent nonlinear coupling. Prominent techniques rely on studying linear perturbations in the presence of a strong magnetic field, for example in measuring the attenuated total reflection [5], or positive longitudinal magnetoconductance [23]. Piezomagnetic probes have also been used to identify the dynamical bulk orbital magnetization induced by strain [24], while others relied on detecting surface plasmons [25, 5]. These approaches have proven valuable for detecting equilibrium responses, however, collective dynamics often rely on nonlinearities that induce a more complex evolution.
Figure 1: Schematic of the general excitation and detection set-up (a) Excitation scheme: two incident linearly polarized beams, $A^{(1)}(\omega_{1},\bm{q}_{1})$ and $A^{(1)}(\omega_{2},\bm{q}_{2})$, come in at a relative angle of $\pi-2\varphi$, with respective linear polarizations $\bm{\hat{\varepsilon}}_{1},\bm{\hat{\varepsilon}}_{2}$; frequencies $\omega_{1},\omega_{2}$; and momenta $\bm{q}_{1},\bm{q}_{2}$. The two beams are used to induce finite $\bm{E}\cdot\bm{B}$, exciting an axionic collective mode using a multiphoton absorption event. (b) Detection scheme: a linearly polarized probe beam, with polarization $\bm{\hat{\varepsilon}}_{I}$, at normal incidence; upon reflection the beam’s polarization $\bm{\hat{\varepsilon}}_{R}$ will in general be elliptically polarized at a polarization angle that oscillates in phase with the induced axion oscillations.
Here we define an all-optical, contactless protocol to identify nonlinear signatures of dynamical axion excitations in a three-dimensional insulating system, termed a dynamical axion insulator (DAI). Our general approach applies to any system hosting axionic excitations. The protocol has two parts: first, a dynamical axion mode is resonantly excited from a background axion field via two-photon Raman processes. By solving the nonlinear coupling between electromagnetic fields and collective axion modes, we show that the axion’s transient evolution depends sensitively on the incident angle, polarization, and frequencies of the two photons. After excitation, a third probe beam detects the resulting axion electrodynamics by monitoring Kerr angle oscillations, serving as a faithful proxy for the axion field’s dynamics [26, 27]. We apply our protocol to a minimal Hamiltonian model that serves as the basis of describing a variety of candidate materials.
Our all-optical, contactless protocol offers advantages over existing methods. Ahn et al. [27] study static, frequency-dependent axions in magnetic TIs requiring broken time-reversal symmetry. In contrast, our method applies to general axionic materials with fluctuating collective modes. While Li et al. [5] and Gooth et al. [23] infer dynamical axions indirectly via static fields or transport, we directly track axion oscillations using time-resolved Kerr rotation—without linearizing the equations of motion—yielding a more general and sensitive probe.
Axion electrodynamics—We start by considering the axion-modified electromagnetic Lagrangian
$\mathcal{L}_{\rm EM}=\frac{1}{2}\left(\epsilon\bm{E}^{2}-\frac{1}{\mu}\bm{B}^{2}\right)+\frac{\alpha}{\pi}\theta\bm{E}\cdot\bm{B}.$ (1)
The first two terms are the sourceless Maxwell Lagrangian in terms of the electric (magnetic) fields $\bm{E}$ ($\bm{B}$), and $\epsilon$ ($\mu$) is the permittivity (permeability) tensor. The last term proportional to $\bm{E}\cdot\bm{B}$ is the so-called axion response that leads to a magnetoelectric coupling mediated by the $\theta$ term, and $\alpha\approx\frac{1}{137}$ is the fine-structure constant. Note that natural units $c=\hbar=1$ are used throughout unless otherwise specified.
From the axion-modified electromagnetic Lagrangian (1) the current density is given by
$\bm{J}=\frac{\alpha}{\pi}\nabla\theta\times\bm{E}+\frac{\alpha}{\pi}\partial_{t}\theta\bm{B},$ (2)
which quantifies the electromagnetic response of the system due to the axion term. The first term in the above equation is equivalent to an anomalous Hall response proportional to the spatial gradient of the axion field, while the second term is known as the chiral magnetic effect and is only relevant in the presence of time-dependent axion fields.
In a dynamical axion insulator, $\theta$ is a degree of freedom which has its own equations of motion and may depend on space and time, $\theta(\bm{r},t)=\theta_{0}(\bm{r},t)+\delta\theta(\bm{r},t)$, where the first term is the background axion, while the second term $\delta\theta(\bm{r},t)$ is the dynamical axion mode. $\theta_{0}$ is determined from the microscopic band structure alone, whereas $\delta\theta$ is a dynamical field with distinct equations of motion that couple nonlinearly to the electromagnetic fields.
When this is the case, the electromagnetic Lagrangian of Eq. (1) is supplemented by the equations of motion of a spinless bosonic field that characterizes the dynamical axion mode, namely
$\mathcal{L}_{\rm NG}=\frac{\kappa}{2}\left[\left(\partial_{t}\delta\theta\right)^{2}-\Omega_{0}^{2}\delta\theta^{2}\right],$ (3)
where $\kappa$ is a materials-dependent parameter, to be given a physical interpretation as “compressibility” later on when we consider the specific case of an axion insulator. We have also included a finite gap for the collective mode $\sim\Omega_{0}$, which reflects the possibility of a resonance at finite frequency, e.g., due to collective mode pinning from disorder.
As we are dealing with electrodynamics, we assume that the speed at which the axion propagates is orders of magnitude slower than the speed of light; as such the axion is effectively dispersionless provided it has a finite gap $\Omega_{0}$. Comparing to the speed of light, $c\gg\nu$ so that all terms relating to first gradients of the dynamical axion field, $\nabla\delta\theta$, can be safely ignored; these describe effects from axion dispersion through the medium.
The equation of motion for the dynamical axion mode $\delta\theta$ are given by
$\kappa\left(\partial_{t}^{2}+\gamma\partial_{t}+\Omega_{0}^{2}\right)\delta\theta=\frac{\alpha}{\pi}\bm{E}\cdot\bm{B},$ (4)
where we have included a phenomenologically relevant dissipation term $\kappa\gamma\partial_{t}\delta\theta$, with damping rate $\gamma$, which represents the coupling of the order parameter to other degrees of freedom in the bulk. The above equation describes the dynamics of a massive spinless bosonic field, driven by an effective force $\frac{\alpha}{\pi}\bm{E}\cdot\bm{B}$.
[…]
$\delta\theta(q)=\frac{(\alpha/\pi\kappa)\omega_{1}\omega_{2}\mathcal{A}_{1}\mathcal{A}_{2}}{\Omega^{2}+i\Omega\gamma-\Omega_{0}^{2}}\left(\frac{\bm{q}_{1}}{\omega_{1}}-\frac{\bm{q}_{2}}{\omega_{2}}\right)\cdot\left(\bm{\hat{\varepsilon}_{1}}\times\bm{\hat{\varepsilon}_{2}}\right),$ (5)
where axion frequency is $\Omega=\omega_{1}+\omega_{2}$ and axion momentum $\bm{q}=\bm{q}_{1}+\bm{q}_{2}$ due to energy-momentum conservation. In general, the intensity of the axion response depends on the frequency, momenta, polarization, and angle of the incident beams, see Fig. 1(a).
Importantly, the axion oscillation amplitude in Eq. (5) depends on: (i) a geometric volume set by polarization vectors $\hat{\bm{\varepsilon}}_{i}$ and momenta $q_{i}/\omega_{i}$, and (ii) a frequency-dependent Lorentzian response. As a second-order nonlinear optical process, excitation occurs via two channels: two-photon absorption (sum or second harmonic generation), Fig.2(a), and stimulated Raman scattering (difference frequency generation), Fig.2(b). The channel dependence arises from the parallelepiped volume formed by $\hat{\varepsilon}_{i}$ and $\bm{q}_{1}/\omega_{1}-\bm{q}_{2}/\omega_{2}$, as shown in Fig.2(c), constraining the response by beam geometry.
The efficacy of this two-photon excitation protocol for a variety of different incident beam angles $\varphi$ and frequencies $\omega_{i}$ in the maximally-crossed polarization geometry, i.e., we keep the polarizations of the excitation beams orthogonal $\hat{\varepsilon}_{1}\cdot\hat{\varepsilon}_{2}=0$, c.f., Fig. 1. In Fig. 3, we show the induced dynamical axion amplitude $|\delta\theta|^{2}$ as a function of incident photon frequencies $\omega_{1},\omega_{2}$ for different incident angles, $\varphi=0,\pi/4$, and $\pi/2$. Fig. 3 clearly shows how varying degrees of freedom of the incident beam can excite the axion field, with the inelastic stimulated Raman channel for collinear beams resulting in the strongest response.
Figure 2: Schematic depiction of different axion excitation pathways. (a) Two photons can be absorbed in order to produce an axion excitation, which is most efficient when $\omega_{1},\omega_{2}\sim\Omega_{0}/2$. (b) Stimulated Raman can excite the axion when an incoming photon inelastically scatters and emits an axion collective mode, leaving the sample with a different energy. (c) Parallelpiped volume generated by triple product in Eq. (5).
Figure 3: This figure illustrates two distinct axion excitation pathways: two-photon absorption (upper half-plane) and stimulated Raman scattering (lower half-plane), with excitation amplitude $|\delta\theta|^{2}$ shown via color bars. The amplitude is plotted against normalized beam frequencies $\omega_{1}/\Omega_{0}$ and $\omega_{2}/\Omega_{0}$, with $\Omega_{0}=1$, and beam geometries indicated atop each subfigure. (a) For $\varphi=0$ (anti-parallel beams), two-photon absorption dominates. (b) At $\varphi=\pi/2$ (parallel incidence), two-photon absorption is suppressed, and Raman excitation is maximal. (c) For $\varphi=\pi/4$ (orthogonal beams), both channels are present, with Raman excitation strongly enhanced. The grey line at $y=0$ separates the two excitation regimes.
Axion detection—We now move on to demonstrate how the axion collective mode may be detected. The key idea is to use a linearly polarized detection beam of known frequency at normal incidence, which when reflected off the surface of the material will posses a time-dependent Kerr effect induced by the presence of the dynamical axion  [26].
The proposed pump-probe style protocol is laid out in Fig. 1, where the pump beam excites the dynamical axion mode in the bulk and is subsequently probed through the vacuum by the reflected field. The background axion in the material is assumed to break both inversion and time-reversal symmetry, such that it gives rise to a bulk Hall conductivity $\nabla\theta_{0}$, and bulk magnetic conductivity $\partial_{t}{\theta}_{0}$, respectively [c.f. Eq. (2)]. We note that in the presence of time-reversal symmetry the Hall effect, related to $\nabla\theta_{0}$, will vanish; correspondingly, in the presence of inversion symmetry the chiral magnetic effect, related to $\partial_{t}\theta_{0}$, will vanish.
The dispersion relation of electromagnetic fields in the material is obtained by solving the equations of motion of the electromagnetic field $\epsilon\partial^{2}_{t}{\bm{A}}-\frac{1}{\mu}\nabla^{2}\bm{A}=\partial_{t}\bm{J}$, where $\epsilon$ is the relative permittivity, $\mu$ is the relative permeability, and $\bm{J}$ is the current given in Eq. (2). Without loss of generality we assume a non-magnetic, spatially isotropic medium such that the permeability and permittivity tensors are simplified to $\mu^{-1}=\mathds{1}$ and $\epsilon=\epsilon\mathds{1}$, where $\epsilon$ is a constant characterizing the material. Furthermore, we consider propagation along the z-axis and solve the equations of motion in the plane orthogonal to the direction of propagation. The refractive index given by (see S.I.) $n_{\pm}=\left(\frac{\epsilon\pm\epsilon^{\prime}}{1\pm\mu^{\prime}}\right)^{1/2}$, with corresponding dispersion $k=\omega n_{\pm}$, where $\epsilon^{\prime}=\omega^{-1}{\partial_{z}\theta}$ is related to the electric conductivity and $\mu^{\prime}=k^{-1}{\partial_{t}\theta}$ is related to the magnetic conductivity.
The reflectance and transmittance amplitudes can be obtained via the continuity of electric and total magnetic fields at the interface, given as $\bm{E}^{I}+\bm{E}^{R}=\bm{E}^{T}$, and $\bm{B}^{I}+\bm{B}^{R}=\bm{B}^{T}+\frac{\alpha}{\pi}\delta\theta\bm{E}^{T}$ (see S.I.). Without loss of generality, here we only include the dynamical axion $\delta\theta$ contributions to the Fresnel equations, and set the background axion at the boundary to zero. Any such contributions, as in $\mathds{Z}_{2}$ topological insulators where the axion takes quantized values, can be incorporated in the equilibrium position of the dynamical axion mode $\delta\theta$. Using the matching conditions, the coefficients of reflection are computed as
$R_{\pm}=\pm\frac{1-(n_{\pm}\mp i\frac{\alpha}{\pi}\delta\theta)}{1+(n_{\pm}\mp i\frac{\alpha}{\pi}\delta\theta)},$ (6)
where $R_{\pm}=\frac{A_{\pm}^{R}}{A_{\pm}^{I}}$ is the ratio between the reflected and initial amplitudes of the fields. The reflection coefficients are, in general, different depending on the refractive index of left- and right-circularly polarized light, and the dynamical axion angle.
By linearizing Eq. (6) for small $\delta\theta$ one obtains the following expression for the polar Kerr rotation angle:
$\Theta_{K}\approx\frac{\alpha}{\pi}\frac{(-2+n_{+}^{2}+n_{-}^{2})}{(n_{+}^{2}-1)(n_{-}^{2}-1)}\delta\theta.$ (7)
The final result of this work is contained within the above equation, which shows the modulation of the reflected beam’s polar Kerr rotation angle and its dependence on the collective fluctuations of the axion field, c.f., Eq. (5). This serves as our smoking gun optical signature to unambiguously deduce the presence of a dynamical axion field in the material.
The transient dynamics of the polar Kerr rotation $\Theta_{K}$ approximately tracking the dynamical axion field oscillations is depicted in Fig. 4. We emphasize that the amplitudes shown are not to scale, rather this figure is meant to illustrate the coherence between $\Theta_{K}$ and $\delta\theta$. The axion equation of motion Eq. (4) can be understood as a driven, damped harmonic oscillator, whereby following excitation the damping term $\gamma$ causes a ring down of the axion field. To simplify the dynamics, we take an effective force given by a Gaussian wavepacket $\bm{E}\cdot\bm{B}\propto e^{(t/\sigma)^{2}}cos(\Omega_{D}t)$, where $\sigma$ is the spread, and $\Omega_{D}$ is the driving frequency. As a result, the dynamical axion mode and polar Kerr rotation angle oscillate in-phase at the frequency of the drive, with the amplitude maximized when on-resonance, i.e., when $\Omega_{D}=\Omega_{0}$.
Figure 4: Transient Kerr rotation reveals signatures of excited axion modes. Using a pump profile $\bm{E}\cdot\bm{B}\propto e^{(t/\sigma)^{2}}\cos(\Omega_{D}t)$ tuned to two-photon resonance ($\Omega_{D}=\Omega_{0}$), we drive out-of-phase axion oscillations $\delta\theta(t)$ with relative phase $\pi/2$ set by damping $\gamma$. These oscillations are reflected in the time-dependent Kerr angle $\Theta_{K}(t)$, which closely follows the axion dynamics. Amplitudes are illustrative and not to scale.
Material application—Up to now the treatment has been generic for any three-dimensional axionic material in the presence of strong interactions where the axion becomes a fluctuating field. The minimal model of such a dynamical axion insulator is defined by a set of four, anti-commuting elements, that form the basis of the four-dimensional Clifford algebra $Cl_{4}$, and a chiral operator [28, 10]. The phenomenology of this model is equivalent to a correlated Weyl fermion, where the low-energy Hamiltonian is determined by a three-dimensional linear dispersion and a mass with two degrees of freedom, e.g., amplitude and phase. The topology of the ground state at half-filling is characterized by an axion field, which can even become quantized when chiral symmetry is preserved. Within this model, the axion field is intimately related to strong interactions which can endow a dynamical behavior; specific material realizations are given by the corresponding irreducible representations.
For example, a Weyl semimetal with charge density wave order may host dynamical axion modes. In an uncorrelated Weyl semimetal with no CDW the nodes are gapless and the chirality of the Weyl nodes is given by the topological charge associated to each band crossing, which is identical to the Berry flux $C=+1$ (C=1C=-1) for the right- (left-) handed Weyl particles [29] for a spinful system. However, in the presence of sufficiently strong interactions, a transition to a CDW ordered phase with complex order parameter $\Delta$ can occur.
Due to the chiral anomaly, it is known that the phase of the complex CDW can be identified with the axion response [6] of the material. The total CDW phase $\theta$ is given by $\theta(\bm{r},t)=\bm{Q}\cdot\bm{r}+Q_{0}t+\delta\theta(\bm{r},t)\,$, where the first two terms are dubbed the background axion field $\theta_{0}=\bm{Q}\cdot\bm{r}+Q_{0}t$ and are determined by the band structure of the material: the separation in momentum-space $\bm{Q}$, and energy $Q_{0}$, between the Weyl nodes. The former is non-zero only in the case of time-reversal symmetry breaking, while the latter is non-zero for inversion-symmetry breaking systems. The final term $\delta\theta(\bm{r},t)$ is the dynamical axion mode, which can be understood as the collective mode associated to the sliding phase mode of the CDW. In this case, the parameter $\kappa$ in Eq. (3) is physically interpreted as the chiral compressibility.
[…]
Conclusion—To conclude, we have shown that, for a generic, three-dimensional, axion insulating system with strong interactions, a two-step excitation-detection protocol may be used to unambiguously detect the presence of a dynamical axion mode in the bulk via its effects on the polar Kerr rotation angle. To model the dynamical axion and determine a novel optical signature we perturbatively solved the nonlinear partial differential equations of Eq. (4) describing its dynamics. We established a two-step, pump-probe style protocol which excites this mode via a two-photon excitation scheme, and subsequently detects it with a third probe beam. We showed that the intensity of the axion excitation is stronger for parallel collinear beams relative to other beam configurations, understood as the stimulated Raman scattering pathway. Our key result is the proposed smoking-gun optical signature, manifested in the modulation of the polar Kerr rotation which approximately tracks the oscillations in the underlying axion field.
Definitive excitation of the dynamical axion enables control of its nonlinear magnetoelectric effects for technological use. This could advance optical technologies such as nonreciprocal thermal emitters  [33, 34], rectifiers, and infrared upconversion detectors  [35, 36, 37], without relying on strong magnetic fields. In photonics, the axion’s strong electric-magnetic coupling may enable quantum logic gates, long-range quantum communication, and low-power laser systems. Additionally, two-photon excitation offers a pathway to nonlinear quantum spectroscopy and enhanced quantum sensing near the Heisenberg limit.
Acknowledgements.
The authors would like to acknowledge fruitful discussions with Aaron Chou, Fahad Mahmood, Soyeun Kim, Ankit Disa, Yikun Wang, and Emily Been. This work is entirely supported by the Quantum Science Center (QSC), a National Quantum Information Science Research Center of the U.S. Department of Energy (DOE). P.N. gratefully acknowledges support from the Gordon and Betty Moore Foundation grant number #8048 and from the John Simon Guggenheim Memorial Foundation (Guggenheim Fellowship).